# ExBody2: Advanced Expressive Humanoid Whole-Body Control

### v1

@@ Appendix A Environments
A.1 Real-world Deployment
@@
Term Expression Weight
DoF position limits $\mathbb{1}(d_{t}\notin[q_{\min},q_{\max}])$ $-10$
DoF acceleration $\|\ddot{d}_{t}\|_{2}^{2}$ $-3e^{-7}$
DoF error $\|d_{t}-d_{0}\|_{2}^{2}$ $-0.5$
Energy $\|\tau_{t}\dot{d}_{t}\|_{2}^{2}$0.001
Linear velocity (z) $\|v^{\text{lin-z}}_{t}\|_{2}^{2}$1
Angular velocity (xy) $\|v^{\text{ang-xy}}_{t}\|_{2}^{2}$0.4
Action rate $\|a_{t}-a_{t-1}\|_{2}^{2}$ $-0.1$
Feet air time $T_{\text{air}}-0.5$ $10$
Feet velocity $\|v_{\text{feet}}\|_{1}$ $-0.1$
Feet contact force $\|F_{\text{feet}}\|_{2}^{2}$ $-0.003$
Stumble $\mathbb{1}(F_{\text{feet}}^{x}>5\times F_{\text{feet}}^{z})$ $-2$
Hip pos error $\|p^{\text{hip}}_{t}-p^{\text{hip}}_{0}\|_{2}^{2}$ $-0.2$
Waist roll pitch error $\|p^{\text{wrp}}_{t}-p^{\text{wrp}}_{0}\|_{2}^{2}$ $-1$
Ankle Action $\|a^{\text{ankle}}_{t}\|_{2}^{2}$ $-0.1$
Table 5: Regularization rewards for preventing undesired behaviors for sim-to-real transfer and refined motion.
@@
DoF position (G1) 23
DoF position (H1) 21
DoF velocity (G1) 23
DoF velocity (H1) 21
Last Action (G1) 23
Last Action (H1) 21
@@
Total Dim (G1) 75*25
Total Dim (H1) 69*25
Table 6: Proprioceptive states used in ExBody2. The rotation information is from IMU. 25 is the length of the history proprioception
@@
Root angular velocity 3
Total dim (G1) 65
Total dim (H1) 63
Table 7: Privileged information used in ExBody2.
@@
Total dim (G1) 69
Total dim (H1) 67
Table 8: Reference information used in ExBody2.
A.2 Simulation Tasks
In this section, we provide detailed state space information for the policy training.
Robot Proprioceptive States The robot proprioceptive states for the teacher and the student policy can be found in Table 6. Note that the student policy is trained on longer history length compared to the teacher, as it cannot observe privileged information but have to learn from a longer sequence of past observations.
Privileged Information The teacher policy leverages privileged information to obtain accurate motion-tracking performance. The complete information about the privileged states is listed in Table 7.
Tracking Target Information Both the teacher policy and student policy also take the motion tracking goal as part of their observations, which consists of the keypoint positions, DoF (joint) positions, as well as root movement information. The detailed components of the motion tracking target can be found in Table 8.
Reward Design In the main paper, we partially introduced our tracking-based reward design. Additionally, our reward also contains other penalties and regularization terms. The regularization reward components and their weights to compute the final rewards are introduced in Table 5. The final reward combines both the regularization with the tracking-based reward to train a robust RL policy.
Action Space The action is the target position of joint proportional derivative (PD) controllers, which is 23 dimensions for Unitree G1 and 21 dimensions for Unitree H1. Note that we are using an upgraded version of the Unitree H1. The older H1 has 19 DOF, while the new H1 has 21 DOF.
@@ Appendix B Model and Training Details
B.1 Baseline Implementation
Exbody[9]: The implementation of Exbody on the Unitree H1 robot is fully consistent with the original Exbody design, tracking only the upper-body keypoints and joint positions. For the Unitree G1 robot, we extended Exbody to tracking the all three dofs of the waist, while keeping other aspects identical. The key differences between Exbody and our method are as follows: Exbody focuses solely on upper-body motion tracking, does not utilize a teacher-student structure, uses a history length of only 5, and performs tracking entirely with local keypoints.
@@ B.1 Baseline Implementation
B.2 CVAE Architecture
Hyperparameter Value
KL Loss Weight $\alpha$ 0.5
Smoothness Loss Weight $\beta$ 0.5
Encoder Attention Layers 4
Decoder Attention Layers 4
Attention Heads 8
Transformer Dimension 512
Latent $z$ Dimension 128
Context Length 50
Motion Prediction Horizon 15
Optimizer AdamW
$\beta_{1},\beta_{2}$ 0.9, 0.999
Learning Rate 3e-4
Warmup Steps 2000
Weight Decay 0.01
Learning Rate Scheduler cosine
Batch Size 512
Total Gradient Steps 60000
Table 9: Hyperparameters related to CVAE’s architecture and training.
Our CVAE model contains a transformer-based encoder and decoder. The encoder leverages bidirectional attention to extract the latent posterior variable $z$ based on past $M-$step motions $m_{t-M:t}$ and future $H$-step motions $m_{t+1:t+H+1}$, and the decoder uses causal attention to predict future $H$-step motions $m_{t+1:t+H+1}$ based on $z$ and $m_{t-M:t}$. Here $m_{t}$ includes the current keypoint positions, root velocity, root angular velocity, position, and rotation.
During training, the transformer encoder $f_{enc}$ takes $m_{t-M:t+H+1}$ and a classification token [CLS] and gets the output sequence. We adopt the corresponding output from the [CLS] token to obtain a posterior Gaussian distribution. A latent variable $z$ is sampled and fed into the decoder alongside $m_{t-M:t+H+1}$ to generate $m_{t+1:t+H+1}$. During inference, $z$ is set to the mean of the prior (i.e. zero). The CVAE is trained on the AMASS CMU motion capture dataset, conditioned on the past 50 frames of motions to predict the future 15 motions. We leverage a similar architecture and temporal ensembling strategy as proposed in [79] to obtain the predicted future motion. We provide the training and architecture hyper-parameters of CVAE in Table 9. During deployment, the humanoid robot takes the predicted motion from the CVAE and tracks it seamlessly in the real world.
Figure 5: Sim-to-real experiment results showcasing striding walk motions across SMPL, simulation, and real-world environments.
B.3 Policy Training Hyper-parameters
ExBody2 adopts a teacher-student training framework. The teacher policy is trained with standard PPO [60] algorithm on privileged information, tracking target and proprioceptive states. The student policy is trained with Dagger [59] without privileged information, but using longer history. For both teacher and student policies, we concatenate the corresponding inputs and feed them into MLP layers for policy learning. We provide the detailed training hyper-parameters for our teacher and student policy in Table 10.
@@ B.3 Policy Training Hyper-parameters
Table 10: Hyperparameters related to the teacher and student policy’s training.
Appendix C More results of ExBody2
Figure 4 and Figure 5 provide illustrative examples of ExBody2’s motion tracking quality. Figure 4 highlights representative correspondences during specific motion tracking processes, showcasing the alignment between the original human SMPL model, simulation results, and real-world tracking outputs, which demonstrate high consistency. Additionally, Figure 5 presents a complete motion sequence, illustrating the correspondences as the motion progresses over time.

### v2

@@ Appendix A Environments
Figure 6: Illustration of ExBody2’s multi-source application, demonstrating how VR, RGB, motion capture, and generative models can be combined to produce diverse humanoid behaviors. (a) Motion Datasets: specialized policies (e.g., kung fu, dancing) finetuned on specialist motion datasets. (b) Real-time Whole-body Mimic: real-time replication of human motions from monocular RGB via HybrIK. (c) Motion Synthesis: a CVAE-based approach for extended and varied motion generation. Experiments demonstrate ExBody2’s capability to seamlessly integrate multiple motion sources in both simulation and real-world scenarios.
A-A Real-world Deployment
@@
A-B State Space Definition
In this section, we provide detailed information on the state space used for policy training, including proprioceptive states, privileged information, and motion tracking targets.
Robot Proprioceptive States. The robot proprioceptive states for the teacher and the student policy can be found in Table V. Note that the student policy is trained on longer history length compared to the teacher, as it cannot observe privileged information but have to learn from a longer sequence of past observations.
Privileged Information. The teacher policy leverages privileged information to obtain accurate motion-tracking performance. The complete information about the privileged states is listed in Table VI.
Tracking Target Information. Both the teacher policy and student policy also take the motion tracking goal as part of their observations, which consists of the keypoint positions, DoF (joint) positions, as well as root movement information. The detailed components of the motion tracking target can be found in Table VII.
Action Space. The action is the target position of joint proportional derivative (PD) controllers, which is 23 dimensions for Unitree G1.
@@ A-B State Space Definition
DoF position 23
DoF velocity 23
Last Action 23
@@
Total Dim 75*10
TABLE V: Proprioceptive states used in Exbody2. The rotation information is from IMU. 10 is the length of the history proprioception
@@
Total dim 62
TABLE VI: Privileged information used in Exbody2.
@@
Total dim 69
TABLE VII: Reference information used in Exbody2.
@@ Appendix B Model and Training Details
B-A Baseline Implementation
Exbody [4]: The implementation of Exbody is consistent with the original Exbody design, tracking only the upper-body keypoints and joint positions. For the Unitree G1 robot, we extended Exbody to tracking the all three dofs of the waist, while keeping other aspects identical. The key differences between Exbody and our method are as follows: Exbody focuses solely on upper-body motion tracking, does not utilize a teacher-student structure, uses a history length of only 5, and performs tracking entirely with local keypoints.
@@ B-A Baseline Implementation
B-B Policy Training Hyper-parameters
Exbody2 adopts a teacher-student training framework. The teacher policy is trained with standard PPO [51] algorithm on privileged information, tracking target and proprioceptive states. The student policy is trained with Dagger [50] without privileged information, but using longer history. For both teacher and student policies, we concatenate the corresponding inputs and feed them into MLP layers for policy learning. We provide the detailed training hyper-parameters for our teacher and student policy in Table VIII.
@@ B-B Policy Training Hyper-parameters
TABLE VIII: Hyperparameters related to the teacher and student policy’s training.
B-C Reward Design
In the main paper, we partially introduced our tracking-based reward design. Additionally, our reward also contains other penalties and regularization terms. The regularization reward components and their weights to compute the final rewards are introduced in Table IX. The final reward combines both the regularization with the tracking-based reward to train a robust RL policy.
Term Expression Weight
DoF position limits $\mathbb{1}(d_{t}\notin[q_{\min},q_{\max}])$ $-10$
DoF acceleration $\|\ddot{d}_{t}\|_{2}^{2}$ $-3e^{-7}$
DoF error $\|d_{t}-d_{0}\|_{2}^{2}$ $-0.1$
Action rate $\|a_{t}-a_{t-1}\|_{2}^{2}$ $-0.1$
Feet air time $T_{\text{air}}-0.5$ $10$
Feet contact force $\|F_{\text{feet}}\|_{2}^{2}$ $-0.003$
Stumble $\mathbb{1}(F_{\text{feet}}^{x}>5\times F_{\text{feet}}^{z})$ $-2$
Waist roll pitch error $\|p^{\text{wrp}}_{t}-p^{\text{wrp}}_{0}\|_{2}^{2}$ $-0.5$
Ankle Action $\|a^{\text{ankle}}_{t}\|_{2}^{2}$ $-0.1$
TABLE IX: Regularization rewards for preventing undesired behaviors for sim-to-real transfer and refined motion.
Metrics
(a) Eval. on $\mathcal{D}_{50}$
$\mathcal{D}_{50}$ 0.1375 0.0627 0.0571 0.0682 0.0753 0.0626 0.0928
$\mathcal{D}_{250}$ 0.1454 0.0669 0.0600 0.0738 0.0870 0.0689 0.1119
(b) Eval. on $\mathcal{D}_{\text{CMU}}$
$\mathcal{D}_{50}$ 0.3509 0.1076 0.1074 0.1076 0.1338 0.1285 0.1410
$\mathcal{D}_{250}$ 0.2834 0.1048 0.1021 0.1073 0.1148 0.1012 0.1335
$\mathcal{D}_{\text{CMU}}$ 0.2622 0.1071 0.1036 0.1110 0.1291 0.1129 0.1512
(c) Eval. on $\mathcal{D}_{\text{ACCAD}}$
$\mathcal{D}_{50}$ 0.4226 0.1277 0.1210 0.1330 0.1720 0.1618 0.1861
$\mathcal{D}_{250}$ 0.3533 0.1234 0.1141 0.1315 0.1421 0.1223 0.1692
TABLE X: Dataset Ablation Study: Evaluation on $\mathcal{D}_{50}$, $\mathcal{D}_{\text{CMU}}$, and $\mathcal{D}_{\text{ACCAD}}$ datasets with models trained on various datasets. Statistically significant results are highlighted in bold across 5 random seeds.
Appendix C Empirical Analysis of Dataset Selection
In the main paper, we propose a Feasibility–Diversity Principle, which posits that a good motion dataset for humanoid tracking must be:
  1. 1.
    Diverse enough (especially in upper-body movements) to ensure the learned policy can generalize beyond very simple or repetitive actions.
  2. 2.
    Feasible enough that lower-body motions do not exceed the robot’s mechanical limits, avoiding extreme samples (e.g., tumbling, handstands) that hamper training.
To illustrate how we arrived at this principle, We manually design three datasets of varying sizes, where the largest being the complete CMU dataset. The remaining datasets, with sizes 50, and 250, are subsets of the CMU dataset, each constructed with different levels of action diversity:
  • 50-action dataset ($\mathcal{D}_{50}$): A minimal set containing only fundamental and mostly static actions (e.g., standing, simple walking). While highly feasible, it lacks diversity in both upper and lower limb motions.
  • 250-action dataset ($\mathcal{D}_{250}$): A moderate-sized set extending $\mathcal{D}_{50}$ with additional upper-limb variations (e.g., arm gestures, some dance moves) and moderately dynamic lower-body actions (e.g., running, mild jumps). Crucially, it avoids highly extreme motions that are difficult for the robot to replicate.
  • Full CMU dataset ($\mathcal{D}_{CMU}$): The complete CMU motion-capture repository of 1,919 sequences, including extreme movements like push-ups, rolling on the ground, and somersaults. Although highly diverse, it contains many infeasible actions that can introduce significant training noise.
We train separate policies with our Exbody2 framework on each dataset above and test them on three different evaluation sets:
  1. 1.
    $\mathcal{D}_{50}$ (in-distribution for the simplest data).
  2. 2.
    $\mathcal{D}_{CMU}$ (the full, more complex dataset).
  3. 3.
    $\mathcal{D}_{ACCAD}$, an out-of-distribution set containing actions not found in any of the training subsets.
Table X summarizes our findings:
  • Evaluation on $\mathcal{D}_{50}$: Policies trained on $\mathcal{D}_{50}$ unsurprisingly achieve the highest tracking accuracy for in-distribution actions, as reflected in metrics across all categories. This suggests that additional data does not necessarily benefit in-distribution tasks. While the policy trained on $\mathcal{D}_{250}$ performs similarly to $\mathcal{D}_{50}$, the policy trained on $\mathcal{D}_{CMU}$ exhibits a substantial drop in tracking accuracy.
  • Evaluation on $\mathcal{D}_{CMU}$: Policies trained on $\mathcal{D}_{250}$ achieve the best performance on $\mathcal{D}_{CMU}$, surpassing those trained on the full $\mathcal{D}_{CMU}$ dataset. Due to the limited diversity of the $\mathcal{D}_{50}$ dataset, especially in upper limb movements, the $\mathcal{D}_{50}$-trained policy struggles to maintain high accuracy for out-of-distribution actions. Unexpectedly, the $\mathcal{D}_{250}$-trained policy generalizes better than the one trained on $\mathcal{D}_{CMU}$. This result underscores that noisy datasets degrade policy performance, as the policy may expend unnecessary effort on tracking infeasible actions, lowering the accuracy of feasible actions.
  • Evaluation on $\mathcal{D}_{ACCAD}$: This experiment further emphasizes the importance of clean datasets. Here, the ACCAD dataset ($\mathcal{D}_{ACCAD}$) consists of actions that are entirely not in the training data. The policy trained on $\mathcal{D}_{250}$ outperforms the others, achieving the best tracking accuracy. Additionally, the $\mathcal{D}_{250}$ and $\mathcal{D}_{CMU}$-trained policies perform relatively well in velocity tracking. However, the $\mathcal{D}_{50}$-trained policy suffers from substantial tracking errors, suggesting the limitations of a small, simple dataset in handling unseen data.
In conclusion, these results validate the core insight behind our Feasibility–Diversity Principle. A small dataset ($\mathcal{D}_{50}$) is indeed easy for the policy to master but lacks sufficient variety to generalize well. On the other hand, a fully unfiltered large dataset ($\mathcal{D}_{CMU}$)—while highly diverse—contains many motions well beyond the robot’s capabilities, introducing detrimental noise. The $\mathcal{D}_{250}$ subset thus provides the best balance between feasible lower-body motions and diverse upper-body actions, enabling our policy to learn robust and expressive whole-body control.
Appendix D Policy Ablation and Additional Results
(a) History Length Ablation
Exbody2-History10 (Ours) 0.2930 0.1000 0.0960 0.1040 0.1079 0.0953 0.1253
Exbody2-History0 0.4151 0.1047 0.1010 0.1081 0.1119 0.0986 0.1303
Exbody2-History25 0.2950 0.1032 0.0984 0.1078 0.1128 0.0965 0.1351
Exbody2-History50 0.2648 0.1004 0.0956 0.1051 0.1114 0.0967 0.1317
Exbody2-History100 0.3242 0.1063 0.1001 0.1122 0.1225 0.1050 0.1466
(b) DAgger Ablation
Exbody2(Ours) 0.2930 0.1000 0.0960 0.1040 0.1079 0.0953 0.1253
Exbody2-w/o-DAgger 0.4195 0.1150 0.1106 0.1198 0.1496 0.1416 0.1607
TABLE XI: Self Ablation Study: Evaluation of different configurations of our method on dataset $\mathcal{D}_{CMU}$. The table is divided into two parts: (a) History Length Ablation and (b) DAgger Ablation.
D-A Ablation on Policy Training
We conduct ablation studies on our policy design to highlight the effectiveness of both (i) the history length for the student policy and (ii) the teacher–student (DAgger) distillation.
History length.
We test student policies trained with different history lengths in Table XI (a). When no extra history is used, the policy struggles to learn effectively. Among the non-zero history lengths, most policies perform similarly while the history length of 10 yields the best results, which is used by us in the main experiments. Longer history lengths increase the difficulty of fitting the privileged information, ultimately reducing tracking performance.
Teacher–student distillation.
Table XI (b) shows that removing DAgger-style distillation severely degrades performance. Without privileged velocity guidance, the student policy must learn velocity tracking directly from raw observations, making it harder to track fast or dynamic motions accurately.
D-B Distribution-Guided Threshold Selection
To choose filtering thresholds in a principled manner, we first analyze the error distribution of the base policy across the entire dataset. Figure 8 presents the empirical cumulative distribution of $e(s)$, with the x-axis indicating the percentile of motion sequences (from lowest to highest error) and the y-axis displaying the corresponding error value.
[…]
Figure 8: Empirical CDF of the base policy’s error metric $e(s)$ on the entire $\mathcal{D}_{\mathrm{CMU}}$ dataset. The horizontal axis indicates the percentile of motion sequences from $0\%$ (lowest error) to $100\%$ (highest error), while the vertical axis shows $e(s)$. We overlay dashed horizontal lines at key thresholds ($\tau=0.075,0.10,0.125,0.15,0.175$) to illustrate how we systematically determine feasible versus unfeasible motions based on the empirical distribution.
D-C Real-world Results Visualization
Figure 7 illustrates how ExBody2 successfully replicates various motions in both simulation and real-world settings. We align each frame’s pose from (i) the reference SMPL animation, (ii) our simulated humanoid robot, and (iii) the real robot deployment. These snapshots confirm that our learned policy retains high fidelity to the target motion, including lower-body poses critical for balance. Additional results can be viewed in the supplementary video.
Appendix E ExBody2’s Multi-source Demonstration
One key advantage of ExBody2 is its flexibility in handling multiple motion sources. In the main text, we focus primarily on motion capture data (i.e., offline datasets). Below, we highlight two other sources—Real-time Whole-body Mimic (RGB-based) and Motion Synthesis (latent generative model)—that can drive ExBody2 for more interactive and long-horizon tasks. Figure 6 visually summarizes these capabilities alongside possible VR or IMU-based streams.
E-A Real-time Whole-body Mimic (RGB Input)
We implement a real-time tracking pipeline that uses only monocular RGB input to mimic human movements. Our system first applies the HybrIK algorithm [30] to extract 3D human poses from each image frame. We then retarget this sequence of poses to the robot’s kinematic structure and feed it into the ExBody2 whole-body policy. Because our policy is trained to be robust to partial or noisy signals, it can accommodate real-time streaming of 3D keypoints and still maintain stable lower-body tracking. Figure 6 (b) demonstrates a user controlling the robot to lift and carry an object, showcasing responsive teleoperation.
Relying on monocular pose estimates is more lightweight than requiring a full-body Mocap or multi-camera setup. Although the 3D pose can be less accurate than multi-view solutions, our control policy’s robust design helps it remain stable even under potential keypoint noise.
E-B Motion Synthesis for Extended Behaviors
We further incorporate a Conditional Variational Autoencoder (CVAE) to generate new motion segments based on a short sequence of past motions, as Figure 6 (c) illustrated. During inference, each latent code $z$ is sampled (or set to the prior mean) to produce new motion trajectories that seamlessly continue from the current pose. Unlike naive random sampling, the CVAE ensures continuity by conditioning on past pose context and penalizing abrupt transitions with a smoothness loss.
Training details.
The CVAE is trained on a broad set of humanoid motion clips, optimizing a reconstruction loss plus KL-divergence for the latent space. We also add a small penalty for high-frequency velocity changes, improving the realism of the generated motions.
Integration with ExBody2.
The generated motion frames are retargeted in exactly the same way as a regular Mocap clip, so the policy sees no difference. This allows the robot to perform extended, varied sequences—e.g., spontaneously transitioning from walking to an upper-body gesture—without needing to rely on a fixed database of motion capture clips.